A proton is passing the origin. The magnetic field at the (x,y,z) position ( 1 mm , 0 mm, 0 mm) is 1.0×10−13ȷ^T. The field at (0 mm, 1 mm , 0 mm) is −1.0×10−13ı^T.

What is the speed of the proton?
Answer #1

6.25*10^-25 m/s

k J 41 1013 10 15 10 3 m r= 1x10 3 lo= 4TX 107 1x10-13; = 10781.6810-19 x (W x 7) V=Vai + Vy j & Vzk - 1013 v = (-14 x 103 E-